If the K_b of a weak base is 1.4 times 10^-6, what
Answer #1

Let the base be BOH

Thus, BOH(aq) <------> B+(aq) + OH-(aq)

Initially, [BOH] = 0.32 M & [B+] = [OH-] = 0 M

Let at eqb., [BOH] = (0.32-x) M & [B+] = [OH-] = x M

Thus, Keq = x2/(0.32-x)

or, 1.4*10-6 = x2/(0.32-x)

or, x = 6.68*10-4 = [OH-]

Thus, pOH = -log[OH-] = 3.175

pH = 14 - pOH = 10.825